OpenStudy (anonymous):

50=2^t, i forget what to do to get the t out of exponent.

OpenStudy (anonymous):

u need to take log on both sides

OpenStudy (anonymous):

just realized that,. lol

jimthompson5910 (jim_thompson5910):

apply logs to both sides to get 50 = 2^t log(50) = log(2^t) log(50) = t*log(2) ... use the rule log(x^y) = y*log(x) now you can isolate t